OpenStudy (osanseviero):

I have a system of 3 equations and five unknown vatiables. Can I solve it with Gauss method?

OpenStudy (anonymous):

like bounded area ??

OpenStudy (osanseviero):

I have this: 3x+3y+4z-5w+3v=5 x+y+2w-v=2 y+2z-w+2v=0

OpenStudy (mathmale):

With only three equations in five unknowns, no, you cannot completely solve this sytem. However, through using the Gaussian elimination method, you can eliminate some of the variables and end up with perhaps two equations in the remaining variables. That's all.

OpenStudy (anonymous):

by gauss methaod you eliminate one by one the variables from the equations. Sincé you have 5 variables to eliminate and only 3 equations, you will no be able to make it

OpenStudy (osanseviero):

Ok, thanks :)